# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v2

@@ 3 A two-stage league-style algorithm
This section introduces a two-stage league-style algorithm that also enjoys a provable scaling law, but under technical assumptions that are different from those required by the knockout-style algorithm.
Motivation: relax Assumption 1.
Recall that Assumption 1 requires that $p_{\text{comp}}>0.5$ holds true for any pair of correct and incorrect solutions, which renders it somewhat restricted and non-robust. This limitation motivates us to propose a league-style algorithm that, as we will show very soon, enjoys a provable scaling law under certain assumptions about the average win rate of a correct candidate solution against a distribution of opponents, rather than against a single one.
@@
where $\phi(r_{i}(j),y_{i},y_{o_{i}(j)})$ denotes the score assigned, based on the response $r_{i}(j)$, to the candidate $y_{i}$ in its comparison against the opponent $y_{o_{i}(j)}$, e.g., 1 for a win, 0 for a loss, and 0.5 for a tie. Finally, the candidate with the highest average win rate $\widehat{\mu}_{i}$ is chosen (with ties broken randomly) as the output of the algorithm.
Regarding computational efficiency, the proposed algorithm (with hyperparameters $N$ and $K$) requires $N$ fully parallelizable LLM calls for the generation stage, and $N\times K$ fully parallelizable LLM calls for the aggregation stage.
@@
With this notation in place, we present our key assumption below.
@@
For the input problem $x$, there exist $p_{\text{gen}}>0$, $\Delta>0$, and a way of dividing the set $\mathcal{Y}$ of all possible solutions into three disjoint subsets $\mathcal{Y}=\mathcal{Y}_{\text{cs}}\cup\mathcal{Y}_{\text{cw}}\cup\mathcal{Y}_{\text{inc}}$ (where “cs”, “cw” and “inc” in the subscripts stand for “correct-and-strong”, “correct-but-weak” and “incorrect”, respectively), such that
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\in\mathcal{Y}_{\text{cs}})\geq p_{\text{gen}}>0\quad\text{and}\quad\min_{y\in\mathcal{Y}_{\text{cs}}}\mu_{y}-\max_{y\in\mathcal{Y}_{\text{inc}}}\mu_{y}\geq\Delta>0.$
In other words, we assume that the LLM can generate a correct-and-strong candidate solution with non-zero probability, whose average win rate against $\mathcal{M}_{\text{gen}}$ is higher (by a strictly positive margin) than that of any incorrect solution. Note that Assumption 2 can be tolerant of systematic errors made by LLMs in comparing certain pairs of candidates, i.e., it may still hold true when there exist a correct solution $y$ and incorrect solution $y^{\prime}$ such that ${\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime})<0.5$, whereas Assumption 1 fails in such cases.
Intuitively, if Assumption 2 holds true and the hyperparameters $N$ and $K$ are sufficiently large, then with high probability, (i) there exist initial candidates that are correct-and-strong solutions, and (ii) $\widehat{\mu}_{i}$ is an accurate estimate of $\mu_{y_{i}}$ for each $i\in[N]$. These conditions together lead to a correct final output of the algorithm. We formalize this intuition in the following theorem.
@@
If Assumption 2 holds for the input problem, then the failure probability of the league-style algorithm with hyperparameters $N$ and $K$ is bounded by
(final output is ${\mathbb{P}}(\text{final output is incorrect})\leq(1-p_{\text{gen}})^{N}+2Ne^{-{K\Delta^{2}}/{8}}+2Ne^{-{(N-1)\Delta^{2}}/{8}}.$
@@
$N\geq\max\bigg{\{}\frac{1}{p_{\text{gen}}}\log\Big{(}\frac{3}{\delta}\Big{)},\ \frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}+1\bigg{\}}\quad\text{and}\quad K\geq\frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}.$
That is, $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $\max\{1/p_{\text{gen}},1/\Delta^{2}\}$ and $1/\Delta^{2}$ (up to logarithmic factors) respectively.
@@ 4 Experiments
We conduct empirical studies in this section to validate the efficacy of the proposed algorithms and confirm the gains from scaling up their test-time compute, while bridging their practical performance with the theories developed in previous sections. Supplementary empirical results can be found in Appendix C.
@@
We use two challenging datasets for our experiments, namely GPQA [30] and MMLU-Pro [37]. GPQA consists of graduate-level multiple-choice questions that pose a challenge even to human experts with PhD degrees in the corresponding domains. It contains over 1000 questions splitted into three categories (“main”, “diamond” and “extended”), all of which are used in our experiments. MMLU-Pro contains 14 categories of multiple-choice questions, some of which require advanced reasoning for obtaining the correct answers. Due to limited computational resources, we use a randomly sampled subset of 100 questions for each category of MMLU-Pro in our experiments, which leads to a total of 1400 questions; we refer to this subset as MMLU-Pro-S in the rest of this work.
@@
We use two open LLMs in our experiments, namely Llama3.1-70B-Instruct (or Llama3.1 for short) [23] and Qwen2.5-72B-Instruct (or Qwen2.5 for short) [28]. We also consider a Mixed option that uses a mixture of both LLMs [34, 48, 12]: during the generation stage, half of the initial candidates are sampled by Llama3.1 and the other half by Qwen2.5; similarly, when a pair of candidates are compared for multiple times during the aggregation stage, half of them are done by Llama3.1 and the other half by Qwen2.5. The rationale is that the capabilities of different LLMs can be complementary to some extent, and thus using a mixture of them can make it more likely that Assumptions 1 and 2 hold true444 To formalize this intuition, consider a minimal scenario with two LLMs denoted by $M_{1}$ and $M_{2}$, and two problems denoted by $x_{1}$ and $x_{2}$. Suppose that $M_{1}$ is good at solving the first problem $x_{1}$ (with $p_{\text{gen}}=0.2$ and $p_{\text{comp}}=0.7$) but incapable of solving $x_{2}$ (with $p_{\text{gen}}=0$ and $p_{\text{comp}}=0.5$), while the reverse is true for $M_{2}$. The magic of using a mixture of two LLMs is that, both problems now satisfy Assumption 1 with $p_{\text{gen}}=(0+0.2)/2=0.1>0$ and $p_{\text{comp}}=(0.5+0.7)/2=0.6>0.5$ (and thus can be solved by our algorithm), whereas only one problem satisfies the assumption when either LLM is used alone. .
Throughout our experiments, the temperature for LLM decoding is set to 0.5 for the generation stage, and 0.1 for pairwise comparisons during the aggregation stage. Our early exploration suggests that these choices strike a good balance between diversity and preciseness in LLM decoding. During the generation stage, we ask the LLM (via zero-shot CoT prompting [17]) to generate a reasoning process first and then its final answer, i.e., the chosen option for a multiple-choice question. For each pairwise comparison during the aggregation stage, we also leverage zero-shot CoT prompting and ask the LLM to think step by step before deciding which solution in the pair is more plausible, unless specified otherwise. The complete prompt templates for both stages can be found in Tables C.3 and C.3 in Appendix C. To account for the positional bias of LLMs [51, 35], we ensure that when a pair of candidates are compared for multiple times, they are placed in one order within the prompt for half of the comparisons, and in the opposite order for the other half.
@@
Unless specified otherwise, for the knockout-style algorithm, we fix the number of comparisons between each pair of candidates encountered in the knockout tournament at $K=4$; for the league-style algorithm, we consider a round-robin (or “all-play-all” [41]) version of it, where 4 pairwise comparisons are conducted between each of (${N\choose 2}$ pairs of initial candidates. To make the proposed algorithms efficient and scalable in practice, we implement them based on AgentScope [9], a multi-agent framework that supports parallel and distributed computation.
@@
In our experiments, we consider a solution to a multiple-choice question as a correct one if its final answer matches the ground-truth answer, and use accuracy as the performance metric for running a (deterministic or randomized) algorithm once on a dataset, i.e., the proportion of problems that the algorithm solves correctly. This metric is, in expectation, equivalent to the mean success probability of the algorithm on the dataset.
@@ 4.1 Results for the knockout-style algorithm
Figure 2: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm.
In our experiments for the knockout-style algorithm, we scale the number of initial candidates $N$ from 1 to 64. Figure 2 confirms that the accuracy achieved by the knockout-style algorithm improves with $N$ for both datasets and all of Llama3.1, Qwen2.5 and Mixed options. For example, accuracy with the Mixed option shows a $10\%$ improvement on GPQA (from $45\%$ to $55\%$), and a $4.5\%$ improvement on MMLU-Pro-S (from $72\%$ to $76.5\%$). We will see later that improvement in accuracy can be much more significant for certain subsets of either dataset. It is also notable that the Mixed option consistently outperforms Llama3.1 and Qwen2.5, which confirms the previously explained rationales for using a mixture of different LLMs.
@@
Figure 3: The distribution of problems from the GPQA (top) or MMLU-Pro-S (bottom) benchmark, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the empirical results for the knockout-style algorithm using the Llama3.1 (left), Qwen2.5 (middle) or Mixed (right) option. Each plot is annotated with the number of problems satisfying the condition $\widehat{p}_{\text{comp}}>0.5$, $\widehat{p}_{\text{comp}}\leq 0.5$, $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{gen}}=1$. To the right of each plot is a histogram for $\widehat{p}_{\text{comp}}$. Each problem is represented by a circle if it is solved correctly by the knockout-style algorithm with $N=64$, and by a cross otherwise. We neglect problems with $\widehat{p}_{\text{gen}}=0$ or $1$, i.e., problems for which the initial candidate solutions are all incorrect or all correct, since there is no way of obtaining meaningful estimate of $\widehat{p}_{\text{comp}}$ for such problems.
Figure 4: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. The values of accuracy are calculated with new trials of the algorithm, which are statistically independent of the trials used for estimating $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$.
@@
The saturation of accuracy in Figure 2 for large values of $N$ is indeed consistent with our theories. Recall that our theorems developed in Section 2.1 guarantee that the knockout-style algorithm can achieve an arbitrarily high success probability for any input problem that satisfies Assumption 1, namely $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. For a problem that does not, it is still possible that the algorithm has a chance of solving it correctly (since Assumption 1 is a sufficient condition for its success and might not be necessary), but there is no formal guarantee. Consequently, for a benchmark or a distribution of input problems, denoted by $\mathcal{D}$, our algorithm is guaranteed to achieve accuracy (i.e., the average of success probabilities) at least ${\mathbb{P}}_{x\sim\mathcal{D}}(x\text{ satisfies the assumption})$ as its test-time compute scales up.
@@
Motivated by this, we conduct further analysis to enhance our understanding of how the knockout-style algorithm achieves its scaling properties empirically. We start by estimating the parameters $p_{\text{gen}}$ and $p_{\text{comp}}$ in Assumption 1. For each problem from the GPQA and MMLU-Pro-S datasets, we define $\widehat{p}_{\text{gen}}$ as the proportion of the $N=64$ initial candidate solutions with the correct final answer, which serves as a good proxy for $p_{\text{gen}}$. To find a proxy for $p_{\text{comp}}$, we define $\widehat{p}_{\text{comp}}$ by picking all LLM calls for comparing a pair of correct and incorrect solutions throughout the knockout tournament, putting higher weights on the comparison results from later rounds of the tournament, and taking the sum of the weights of all comparisons that identify the right winners [^5] The rationale for weighting the comparison results is explained as follows. In the early rounds of the knockout tournament, the comparison result between a correct-but-weak candidate and an incorrect candidate can cause a negative bias in estimating $p_{\text{comp}}$; similarly, a correct candidate might have a very high win rate against an opponent that is not only incorrect but also very weak, which can cause a positive bias in estimating $p_{\text{comp}}$. In contrast, the correct or incorrect candidates that survive the early rounds of the knockout tournament tend to be stronger ones, which make the comparison results among them (in later rounds of the tournament) more reliable and meaningful for the purpose of estimating $p_{\text{comp}}$. .
@@
Figure 3 characterizes the distribution of problems from the GPQA and MMLU-Pro-S datasets, in terms of $\widehat{p}_{\text{gen}}$ (the X-axis) and $\widehat{p}_{\text{comp}}$ (the Y-axis). On the top half of each scatter plot are the problems with $\widehat{p}_{\text{comp}}>0.5$, most of which are solved correctly by the knockout-style algorithm and represented as circles. These include some problems with small $\widehat{p}_{\text{gen}}$, for which the knockout stage successfully identifies a correct candidate even though the initial candidates are mostly incorrect. We further observe that, compared to Llama3.1 and Qwen2.5, the Mixed option achieves $\widehat{p}_{\text{gen}}>0$ and $\widehat{p}_{\text{comp}}>0.5$ for a larger proportion of problems, which explains its superior accuracy shown in Figure 2. Of course, there also exist problems with $\widehat{p}_{\text{gen}}=0$ or $\widehat{p}_{\text{comp}}\leq 0.5$, which prevent us from boosting the accuracy up to $100\%$ on these benchmarks.
To further consolidate this analysis, we pay special attention to the subset of problems that satisfy $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. These correspond to the conditions stated in Assumption 1, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run a new, independent trial of the knockout-style algorithm (with Llama3.1/Qwen2.5/Mixed) on this subset. The results illustrated in Figure 4 confirm that significant improvements in accuracy (e.g., a 25% increase for Mixed) for such problems can be achieved by scaling up the test-time compute of the algorithm, as predicted by our theories. The algorithm does not achieve 100% accuracy on this subset though, since $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ are only approximate estimate of the true underlying $p_{\text{gen}}$ and $p_{\text{comp}}$ respectively, and thus there still exist problems in this subset that violate Assumption 1.
@@
Interestingly, we observe that for MMLU-Pro-S, the scaling properties of the algorithm vary across different categories of the dataset, and also across the options of which LLM backend(s) to use. For instance, Figure 5 shows that the performance scales well for all of Llama3.1, Qwen2.5 and Mixed in the “engineering” category, while the scaling property of Llama3.1 outperforms the other two options in “philosophy”.
Such results suggest that the gains from scaling up the test-time compute of the knockout-style algorithm depend on the characteristics of both problems and LLM backends. For example, for a reasoning-focused task like “engineering”, LLMs can compare the reasoning processes of two candidate solutions side by side, which provides additional information compared to generating or verifying an individual solution, and thus leads to a large value of $p_{\text{comp}}$ and accurate comparison results. In contrast, for a knowledge-heavy task like “philosophy” that mostly requires memorization of relevant knowledge, one would not expect significant gains from pairwise comparison if the LLM simply does not have the right knowledge embedded within its model weights, in which case $p_{\text{comp}}$ might be close to $0.5$ or worse.
Additional results.
More empirical results for the knockout-style algorithm can be found in Appendix C.1, including the impacts of the hyperparameter $K$ and the use of CoT prompting for the aggregation stage, a comparison with majority voting, and results for each individual category of GPQA and MMLU-Pro-S.

### v3

@@ 3 A two-stage league-style algorithm
Recall that Assumption 1 in the previous section requires $p_{\text{comp}}>0.5$ for any pair of correct and incorrect solutions, which renders it somewhat restricted and non-robust. This motivates us to propose a two-stage league-style algorithm that also enjoys a provable inference scaling law, under technical assumptions analogous to but more robust than those required by the knockout-style algorithm.
@@
where $\phi(r_{i}(j),y_{i},y_{o_{i}(j)})$ denotes the score assigned, based on $r_{i}(j)$, to the candidate $y_{i}$ in its comparison against $y_{o_{i}(j)}$, e.g., 1 for a win, 0 for a loss, and 0.5 for a tie. Finally, the candidate with the highest average win rate $\widehat{\mu}_{i}$ is chosen (with ties broken randomly) as the output of the algorithm.
Regarding computational efficiency, the proposed algorithm requires $N$ fully parallelizable LLM calls for the generation stage, and $N\times K$ fully parallelizable LLM calls for the aggregation stage.
@@
Our key assumption is presented below.
@@
For the input problem $x$, there exist $p_{\text{cs}}>0$, $\Delta>0$, and a way of dividing the set $\mathcal{Y}$ of all possible solutions into three disjoint subsets $\mathcal{Y}=\mathcal{Y}_{\text{cs}}\cup\mathcal{Y}_{\text{cw}}\cup\mathcal{Y}_{\text{inc}}$ (where “cs”, “cw” and “inc” stand for “correct-and-strong”, “correct-but-weak” and “incorrect”, respectively), such that
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\in\mathcal{Y}_{\text{cs}})\geq p_{\text{cs}}>0\quad\text{and}\quad\min_{y\in\mathcal{Y}_{\text{cs}}}\mu_{y}-\max_{y\in\mathcal{Y}_{\text{inc}}}\mu_{y}\geq\Delta>0.$
In other words, we assume that the LLM can generate a correct-and-strong solution with non-zero probability, whose average win rate against $\mathcal{M}_{\text{gen}}$ is higher than that of any incorrect solution. Note that Assumption 2 can be tolerant of systematic errors by LLMs in comparing certain pairs of candidates, i.e., it may still hold true when there exist a correct solution $y$ and incorrect solution $y^{\prime}$ such that ${\mathbb{E}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}\phi(r,y,y^{\prime})<0.5$, whereas Assumption 1 fails in such cases.
Intuitively, if Assumption 2 holds true and the hyperparameters $N$ and $K$ are sufficiently large, then with high probability, (1) there exist initial candidates that are correct-and-strong solutions, and (2) $\widehat{\mu}_{i}$ is an accurate estimate of $\mu_{y_{i}}$ for each $i\in[N]$. These conditions together lead to a correct final output of the algorithm. We formalize this intuition in the following theorem.
@@
If Assumption 2 holds for the input problem, then the probability that the league-style algorithm (with hyperparameters $N$ and $K$) returns an incorrect final output is bounded by
${\mathbb{P}}(\text{failure})\leq(1-p_{\text{cs}})^{N}+2Ne^{-{K\Delta^{2}}/{8}}+2Ne^{-{(N-1)\Delta^{2}}/{8}}.$
@@
$N$ $\geq\max\bigg{\{}\frac{1}{p_{\text{cs}}}\log\Big{(}\frac{3}{\delta}\Big{)},\ \frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}+1\bigg{\}}\quad\text{and}\quad K\geq\frac{8}{\Delta^{2}}\log\Big{(}\frac{6N}{\delta}\Big{)}.$
That is, $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $\max\{1/p_{\text{cs}},1/\Delta^{2}\}$ and $1/\Delta^{2}$ (up to logarithmic factors) respectively.
@@ 4 Experiments
We conduct empirical studies to validate the efficacy and scaling properties of the proposed algorithms, while bridging their practical performance with the theories developed in previous sections.
@@
We use three datasets for our experiments: GPQA [34], MMLU-Pro [43] and MATH-500 [26]. GPQA consists of over 1000 graduate-level multiple-choice questions splitted into three categories (“main”, “diamond” and “extended”), all of which are used in our experiments. MMLU-Pro contains 14 categories of multiple-choice questions, some of which require advanced reasoning while others are more knowledge-focused. Due to limited computational resources, we use a randomly sampled subset of 100 questions for each category of MMLU-Pro in our experiments, which leads to a total of 1400 questions; we refer to this subset as MMLU-Pro-S throughout this work. MATH-500 is a subset of 500 problems from the MATH dataset introduced in [22]. We focus mainly on GPQA in this section, deferring empirical results for MMLU-Pro-S and MATH-500 (as well as supplementary results for GPQA) to Appendix C.
@@
We use Llama3.1-70B-Instruct (Llama3.1 for short) [24] and Qwen2.5-72B-Instruct (Qwen2.5 for short) [31] in our experiments, as well as a Mixed option that uses a mixture of both LLMs [40, 54, 13]: during the generation stage, half of the initial candidates are sampled by Llama3.1 and the other half by Qwen2.5; similarly, when a pair of candidates are compared for multiple times during the aggregation stage, half of them are done by Llama3.1 and the other half by Qwen2.5. The rationale is that the capabilities of different LLMs can be complementary to some extent, and thus using a mixture of them can make it more likely that Assumptions 1 and 2 hold true222 To formalize this intuition, consider a minimal scenario with two LLMs denoted by $M_{1}$ and $M_{2}$, and two problems denoted by $x_{1}$ and $x_{2}$. Suppose that $M_{1}$ is effective for the first problem $x_{1}$ (with $p_{\text{gen}}=0.2$ and $p_{\text{comp}}=0.7$) but ineffective for $x_{2}$ (with $p_{\text{gen}}=0$ and $p_{\text{comp}}=0.5$), while the reverse holds true for $M_{2}$. When either LLM is used alone, only one problem satisfies Assumption 1. However, when a mixture of two LLMs is used, both problems now satisfy Assumption 1 with $p_{\text{gen}}=(0+0.2)/2=0.1>0$ and $p_{\text{comp}}=(0.5+0.7)/2=0.6>0.5$, and thus can be solved by our algorithm. . Other models considered in our experiments include QwQ-32B [32], a long-CoT reasoning LLM, and GPT-4o [10], a proprietary API-based LLM; due to high computational or monetary costs, they are tested only on GPQA-diamond for a smaller range of $N$.
@@
We leverage zero-shot chain-of-thought prompting [18] for both generation and aggregation stages of the proposed algorithms. Unless specified otherwise, for the knockout-style algorithm, we fix $K=4$ for Llama3.1/ Qwen2.5/ Mixed, and $K=2$ for GPT-4o/ QwQ-32B; for the league-style algorithm, we consider a round-robin [47] version of it, with $K=4$ comparisons conducted between each of (${N\choose 2}$ pairs of initial candidates. To make the proposed algorithms efficient and scalable in practice, we implement them based on AgentScope [9], a multi-agent framework that supports parallel and distributed computation. Further implementation details can be found in Appendix C.1.
@@
In our experiments, we consider a solution as a correct one if its final answer matches the ground-truth answer, and use accuracy (i.e., the proportion of correctly solved problems) as the performance metric for running a (deterministic or randomized) algorithm once on a dataset. This metric is, in expectation, equivalent to the mean success probability of the algorithm on the dataset.
@@ 4.1 Results for the knockout-style algorithm
Figure 2 confirms that the accuracy of the knockout-style algorithm improves with $N$ for all LLMs on GPQA or GPQA-diamond. For example, the accuracy of Mixed improves by 10 points (from $45\%$ to $55\%$) as $N$ scales to 64, and the accuracy of QwQ-32B improves by 12 points (from $60\%$ to $72\%$) as $N$ scales to 16. We also observe that Mixed consistently outperforms Llama3.1 and Qwen2.5 as $N$ gets larger, which confirms the previously explained rationales for using a mixture of different LLMs.
Comparison with majority voting.
Figure 2 includes results for majority voting, a strong baseline widely adopted in practice. It is observed that, for all LLM backends (except for Llama3.1), the knockout-style algorithm consistently achieves higher accuracy when given the same number $N$ of initial candidates. Caution should be taken here: recall from Section 2.2 that the knockout-style algorithm takes $(K+1)\times N$ LLM calls for solving one problem, i.e., $5\times N$ for Llama3.1/ Qwen2.5/ Mixed and $3\times N$ for GPT-4o/ QwQ-32B, whereas majority voting only requires $N$ LLM calls. Nonetheless, the knockout-style algorithm still has advantage when this is taken into account, e.g., its accuracy at $N=8$ (resp. 4) is higher than that of majority voting at $N=64$ (resp. 16) for Mixed (resp. QwQ-32B). Moreover, based on the trends shown in Figure 2, it is most likely that for majority voting, further increasing $N$ will bring limited performance gains [3] and result in a converged accuracy lower than what can be achieved by the knockout-style algorithm.
Refer to captionRefer to caption
Figure 2: Accuracy versus the number of initial candidates $N$ for the knockout-style algorithm (KO), as well as for majority voting (MV), a strong baseline widely adopted in practice.
@@
The results in Figure 2 are indeed consistent with the theorems developed in Section 2.1, which guarantee that the knockout-style algorithm can achieve an arbitrarily high success probability for any input problem satisfying Assumption 1, namely $p_{\text{gen}}>0$ and $p_{\text{comp}}>0.5$. For a problem that does not, it is still possible that the algorithm has a chance of solving it correctly (since Assumption 1 is a sufficient condition for its success and might not be necessary), but there is no formal guarantee. Consequently, for a benchmark or a distribution of input problems, denoted by $\mathcal{D}$, our algorithm is guaranteed to achieve accuracy at least ${\mathbb{P}}_{x\sim\mathcal{D}}(x\text{ satisfies the assumption})$ as its test-time compute grows. Indeed, if a benchmark contains an extremely difficult problem, e.g., “solve the P versus NP problem”, then any test-time scaling method will fail to achieve 100% accuracy on such a benchmark.
Figure 3: Left: the distribution of GPQA problems, characterized by $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$ that are estimated with the knockout-style algorithm (Mixed). Each problem is represented by a circle if it is solved correctly at $N=64$, and by a cross otherwise. Right: accuracy versus $N$ for the knockout-style algorithm, on a filtered subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. The values of accuracy are calculated with new trials of the algorithm, thus statistically independent of $\widehat{p}_{\text{gen}}$ and $\widehat{p}_{\text{comp}}$.
@@
To further bridge the empirical results with theories, let us start by estimating the parameters $p_{\text{gen}}$ and $p_{\text{comp}}$ in Assumption 1. For each problem, we define $\widehat{p}_{\text{gen}}$ as the proportion of the $N=64$ initial candidate solutions with a correct final answer, which serves as a good proxy for $p_{\text{gen}}$. To find a proxy for $p_{\text{comp}}$, we define $\widehat{p}_{\text{comp}}$ by picking all LLM calls for comparing a pair of correct and incorrect solutions throughout the knockout tournament, putting higher weights on the comparison results from later rounds of the tournament, and taking the sum of the weights of comparisons that identify the right winners [^3] The rationale for weighting the comparison results is explained as follows. In the early rounds of the knockout tournament, the comparison result between a correct-but-weak candidate and an incorrect candidate can cause a negative bias in estimating $p_{\text{comp}}$; similarly, a correct candidate might have a very high win rate against an opponent that is not only incorrect but also very weak, which can cause a positive bias in estimating $p_{\text{comp}}$. In contrast, the correct or incorrect candidates that survive the early rounds of the knockout tournament tend to be stronger ones, which make the comparison results among them (in later rounds of the tournament) more reliable and meaningful for the purpose of estimating $p_{\text{comp}}$. .
@@
Figure 7 in Appendix C.2 characterizes the distribution of GPQA and MMLU-Pro-S problems in terms of $\widehat{p}_{\text{gen}}$ (the X-axis) and $\widehat{p}_{\text{comp}}$ (the Y-axis); one such plot can also be found in Figure 3 (left). On the top half of each scatter plot are problems with $\widehat{p}_{\text{comp}}>0.5$, most of which are solved correctly by the knockout-style algorithm and represented as circles. These include some problems with small $\widehat{p}_{\text{gen}}$, for which the knockout stage successfully identifies a correct candidate even though the initial candidates are mostly incorrect. We further observe that, compared to Llama3.1 and Qwen2.5, the Mixed option achieves $\widehat{p}_{\text{gen}}>0$ and $\widehat{p}_{\text{comp}}>0.5$ for a larger proportion of problems, which explains its superior accuracy shown in Figure 2. To further consolidate this analysis, we pay special attention to the subset of problems satisfying $10<\widehat{p}_{\text{gen}}<$ 1 and $\widehat{p}_{\text{comp}}>0.5$. These are approximations for the conditions stated in Assumption 1, except that those easy problems with $\widehat{p}_{\text{gen}}=1$ are excluded. We run new, independent trials of the knockout-style algorithm on this subset. Figure 3 (right) confirms that significant improvements in accuracy (e.g., 25 points for Mixed) can be achieved for such problems by scaling up $N$, which matches what our theories predict.
@@
Interestingly, we observe that the scaling properties of the algorithm vary across different categories of MMLU-Pro-S, and also across LLM backends. For instance, Figure 4 shows that the performance scales well for all of Llama3.1/ Qwen2.5/ Mixed in the “engineering” category, while the scaling of Llama3.1 outperforms the other two options in “philosophy”. An intuitive explanation is that, for a reasoning-focused task like “engineering”, LLMs can compare the reasoning processes of two candidate solutions side by side, which provides additional information compared to generating or verifying an individual solution, and thus leads to a large value of $p_{\text{comp}}$ and accurate comparison results. In contrast, for a knowledge-heavy task like “philosophy”, one would not expect significant gains from pairwise comparison if the LLM simply does not have the right knowledge embedded within its model weights, in which case $p_{\text{comp}}$ might be close to (or even below) $0.5$.